# Chemical Abundances in the Leiptr Stellar Stream: a Disrupted Ultra-faint Dwarf Galaxy?

### v1

@@ Abstract
Over the course of their lifetimes, stellar systems can be torn apart by the gravitational potential of massive galaxies. Stellar streams form during the intermediate stage of tidal disruption where they preserve their coherence in phase-space before mixing into a stellar halo (Helmi & White, 1999). Current models of galaxy formation predict that there should be a large population of globular cluster and dwarf galaxy stellar streams in the Milky Way’s stellar halo as a result of past accretion events (Helmi, 2020). These systems provide observational evidence to evaluate theoretical models of hierarchical structure formation (e.g., Johnston, 1998; Helmi & White, 1999; Freeman & Bland-Hawthorn, 2002; Bullock & Johnston, 2005) and quantify the amount of dark matter substructure (e.g., Ibata et al., 2002; Johnston et al., 2002; Carlberg, 2009; Varghese et al., 2011; Erkal et al., 2016a). Large photometric surveys such as the Sloan Digital Sky Survey (York et al. 2000) and the Dark Energy Survey (Dark Energy Survey Collaboration et al. 2016) in combination with proper motions from Gaia data releases (Gaia Collaboration et al., 2016; Lindegren et al., 2021; Gaia Collaboration et al., 2023) have revealed over 140 streams (Grillmair & Carlin, 2016; Shipp et al., 2018; Mateu et al., 2018; Mateu, 2023; Ibata et al., 2024; Bonaca & Price-Whelan, 2024).
@@
It is often unclear whether a given stream’s progenitor is a dwarf galaxy, a globular cluster, or possibly a superposition of both (e.g., Carlberg, 2018; Malhan et al., 2021). In some cases, streams are clearly identified with intact progenitors, such as the Sagittarius dwarf galaxy (Ibata et al., 2001; Majewski et al., 2003), the Antlia II and Crater II dwarf spheroidals (Ji et al., 2021; Coppi et al., 2024), and several globular clusters (e.g., Simpson et al., 2020; Hansen et al., 2020a; Gull et al., 2021). The majority of streams do not have associated intact progenitors. Instead, a stream’s morphology and kinematics can provide a strong prior on the progenitor (e.g., Erkal et al., 2016b; Shipp et al., 2018; Ibata et al., 2024). However, a definitive classification requires studying a stream’s chemical abundances (e.g., Li et al., 2019, 2022; Ji et al., 2020a; Martin et al., 2022).
@@
The majority of known streams are thought to have globular cluster progenitors, due to their thin widths and in many cases unresolved metallicity dispersions dex (Li et al., 2022; Ibata et al., 2024). These include the Phoenix (Wan et al., 2020; Casey et al., 2021) and C-19 streams (Martin et al., 2022; Yuan et al., 2022a) which are below the globular cluster metallicity “floor”; the dynamically perturbed GD-1 (Grillmair & Dionatos, 2006; Price-Whelan & Bonaca, 2018) and ATLAS-Aliqa Uma streams (Shipp et al., 2018; Li et al., 2021); disrupted globular clusters embedded within the Indus, Jhelum, and Wukong/LMS-1 dwarf galaxy streams (Hansen et al., 2021; Awad et al., 2024; Limberg et al., 2024); and the 300S stream, which was originally thought to be a dwarf galaxy stream (Simon et al., 2011; Frebel et al., 2013b; Fu et al., 2018) but has since been chemically confirmed to be an accreted globular cluster stream (Li et al., 2022; Usman et al., 2024).
@@
Less common by number are the dwarf galaxy streams. Many of the first discovered streams, such as the Helmi stream (Helmi et al., 1999; Roederer et al., 2010; Limberg et al., 2021a; Matsuno et al., 2022b) and Sagittarius stream (Ibata et al., 2001; Majewski et al., 2003), were easily identified as dwarf galaxies as these are well-populated streams from relatively massive () dwarf galaxy progenitors. More recently, several lower-mass dwarf galaxy streams have also been identified. Based on their chemical patterns, the Indus and Jhelum streams have progenitor mass (Hansen et al., 2021); the Orphan-Chenab stream (Belokurov et al., 2007; Casey et al., 2013; Shipp et al., 2018; Koposov et al., 2019; Erkal et al., 2019) has progenitor mass (Koposov et al., 2019; Hawkins et al., 2023); the Wukong/LMS-1 stream has progenitor mass (Naidu et al., 2020; Yuan et al., 2020; Limberg et al., 2024); the Cetus/Palca stream has progenitor mass (Newberg et al., 2009; Yuan et al., 2022b; Sitnova et al., 2024); and the Elqui stream has progenitor mass (Shipp et al., 2018; Ji et al., 2020a; Li et al., 2022). This does not include fully disrupted and phase-mixed dwarf galaxies like Gaia-Sausage/Enceladus (GSE, Belokurov et al., 2018; Haywood et al., 2018; Helmi et al., 2018; Carrillo et al., 2022), Sequoia (Myeong et al., 2019; Matsuno et al., 2019), Thamnos (Koppelman et al., 2019), and others (Naidu et al., 2020; Helmi, 2020; Deason & Belokurov, 2024).
@@
Current predictions for GC streams are that the Milky Way has globular cluster stellar streams that have yet to be fully phase-mixed (Pearson et al., 2024). In principle, there are also many hundreds of dwarf galaxies that have been tidally disrupted around the Milky Way (e.g., Johnston, 1998; Helmi & White, 1999; Santistevan et al., 2020; Brauer et al., 2022; Gandhi et al., 2024). However, it is harder to identify the dwarf galaxy streams. Current stream searches, especially those relying on Gaia parallaxes (e.g., Ibata et al., 2019, 2024; Dodd et al., 2023), mostly search at nearby galactocentric radii, where globular cluster streams are likely more numerous than dwarf galaxy streams (Johnston et al., 2008; Gómez et al., 2013; Pearson et al., 2024; Bonaca & Price-Whelan, 2024). Dwarf galaxy stars have larger velocity dispersions and metallicity spreads, so at fixed stellar mass they are harder to find compared to globular clusters that are more spatially compact and at single metallicities. Finally, the vast majority of the tidally disrupted dwarf galaxies are very low-mass, with , i.e. ultra-faint dwarf galaxies (e.g., Weisz & Boylan-Kolchin, 2017; Simon, 2019; Brauer et al., 2019; Santistevan et al., 2020; Gandhi et al., 2024), so they should be even harder to identify. Obtaining a full census of the Milky Way’s accretion history thus requires extending stream discoveries down to this regime of ultra-faint dwarf galaxies. Since these ultra-faint dwarf galaxies are also the most dark matter-dominated galaxies that live in the lowest-mass star-forming dark matter halos, understanding the number of these streams in the Milky Way could contribute to understanding reionization and the small scale structure of dark matter (e.g., Bullock & Boylan-Kolchin, 2017; Jethwa et al., 2018; Simon, 2019; Nadler et al., 2021).
@@
In recent years, much progress has been made in finding streams from the lowest-mass dwarf galaxies. The first clear stream around an intact ultra-faint dwarf galaxy was found around the Tucana III system (Drlica-Wagner et al., 2015; Shipp et al., 2018), which has been spectroscopically confirmed (Hansen et al., 2017; Li et al., 2018). More recently, a tidal stream likely emanating from the Hydrus I dwarf galaxy was found (Ibata et al., 2024). Several individual stars with chemistry similar to ultra-faint dwarf galaxy stars have also been found throughout the Milky Way halo (e.g., Casey & Schlaufman, 2017; Roederer, 2017). Motivated by the discovery of the highly -process (rapid neutron-capture process) enhanced ultra-faint dwarf Reticulum II (Ji et al., 2016a; Roederer et al., 2016), several stellar dynamical groups within the Milky Way that are highly enhanced in -process elements have also been suggested to originate from ultra-faint dwarf galaxies (e.g., Roederer et al., 2018; Lee et al., 2019; Limberg et al., 2021b; Gudin et al., 2021), but the probability that the stars in these groups all come from individual ultra-faint dwarf galaxies is statistically unlikely (Brauer et al., 2022). Thus far, there is no progenitor-free bona-fide stellar stream that can be cleanly associated with an ultra-faint dwarf galaxy. The only claimed candidate, Specter (Chandra et al., 2022), has a high mean metallicity that suggests we are only seeing part of the stream from a much higher-mass progenitor dwarf galaxy.
@@
Here we investigate the progenitor of the Leiptr stellar stream, which was previously thought to be a tidally disrupted globular cluster. Leiptr was discovered among a group of inner galaxy streams in the Gaia DR2 catalog by Ibata et al. (2019) using the STREAMFINDER algorithm. They identified candidates from streams with distinct and coherent sky positions, proper motions, and parallaxes that were clearly distinguishable from the Galactic disk or bulge. They determined Leiptr was 48∘ long and located toward the Galactic anticenter with a heliocentric distance of kpc, pericenter distance of kpc, apocenter distance of kpc, and strong retrograde motion of = km s-1 kpc where is the vertical component of angular momentum in a Galactocentric Cartesian frame. Malhan et al. (2022) updated these orbital parameters, estimating the pericenter to be 12.3 0.1 kpc and the apocenter to be 45.1 0.2 kpc, with a resulting eccentricity of 0.57. Leiptr has the highest of stellar streams in their sample, next to Gaia-12. Bonaca et al. (2021) identified Leiptr, Gjöll, GD-1, Phlegethon, Ylgr, and Wambelong as potential members of a highly aligned “plane of streams” that may share a common origin. They suggested the retrograde streams are most likely globular clusters from the Sequoia/Arjuna dwarf galaxy (Myeong et al., 2019; Naidu et al., 2020) based on photometry and clustering in orbital poles; Gjöll in particular has been shown to be connected to the NGC 3201 globular cluster (Hansen et al., 2020a; Riley & Strigari, 2020). Bonaca & Price-Whelan (2024) independently determined that Leiptr has a mean heliocentric distance of 7.4 kpc, mean Galactic radius of 13.8 kpc, angular length of 73∘, angular width of 0.5∘ (74 pc), and a stellar mass of . However, there have been no previous chemical studies of Leiptr.
We present the chemical abundances of four member stars in the Leiptr stream and one non-member using high-resolution spectroscopy. We find that Leiptr is likely the first spectroscopically confirmed ultra-faint dwarf galaxy stream without an intact progenitor. Section 2 explains the observations. Section 3 describes the abundance analysis. Section 4 compares the chemical abundances of Leiptr to those of halo, dwarf galaxy and globular cluster stars. We discuss what these results suggest about Leiptr’s chemical signature and progenitor in Section 5 and conclude in Section 6.
@@ 2 Observations
Of 67 original Leiptr candidate stars from the Gaia DR2 STREAMFINDER catalog (Ibata et al., 2019), we selected five of the brightest candidate stars to observe, giving them arbitrary ID numbers. We obtained spectra of these stars on November 2, 2022 using the 6.5-meter Magellan Clay Telescope’s Magellan Inamori Kyocera Echelle (MIKE; Bernstein et al. 2003) spectrograph at Las Campanas Observatory in Chile. Figure 1 shows the declinations and radial velocities against right ascensions of these stars compared to the other Leiptr candidates from the Gaia DR3 STREAMFINDER catalog in Ibata et al. (2024). One of these five stars, Leiptr-252, is no longer a kinematic Leiptr member star in the Gaia DR3 STREAMFINDER catalog (Ibata et al., 2024), which will also be apparent from its chemistry. This non-member star is shown as a red X in Figure 1 and all subsequent figures.
@@ 3 Abundance Analysis
The chemical abundance analysis was performed using a custom analysis code LESSPayne (Ji et al. in prep), generally following Ji et al. (2020a). As described in Limberg et al. (2024), LESSPayne is a combination of the PayneEchelle (Ting et al., 2019) and SMHR (Casey, 2014) programs. First, we ran a full spectrum fit between Å using a neural network emulator of high-resolution spectra. We did not use the stellar parameters from this fit, but instead used the best-fit spectrum to locate and mask regions where absorption lines were expected to occur and initialize an SMHR analysis. We manually inspected and corrected the normalization for every spectral order before stitching them together. We then fit Gaussian profiles to determine equivalent widths using the line list from Ji et al. (2020a), which is based on atomic data from linemake (Placco et al., 2021), mostly keeping to lines with 4000 Å 7000 Å. Some stronger lines required Voigt profile fits. We also fit some lines using spectral synthesis with the same line lists as Ji et al. (2020a).
@@
We determined chemical abundances using the 1D local thermodynamic equilibrium (LTE) radiative transfer code MOOG (Sneden, 1973) and the ATLAS model atmospheres (Castelli & Kurucz, 2003). Stellar parameters were determined spectroscopically following Frebel et al. (2013a), due to uncertain distances to the Leiptr stars. Effective temperature was derived by balancing Fe I abundance against excitation potential, surface gravity by balancing Fe I and II abundances, and microturbulence by balancing Fe II abundance against reduced equivalent width. Next, we applied the correction to a photometric temperature scale from Frebel et al. (2013a) and redetermined the other stellar parameters. The photometric correction is why we use Fe II for microturbulence, because the Fe I line abundances computed in LTE and with photometric temperatures result in artificially high microturbulence (Ji et al., 2020a), which is possible as we used Fe II lines per star. The corrected effective temperature (), surface gravity (log g), microturbulence (), metallicity ([Fe/H]), and their associated uncertainties (, , , and , respectively) are given in Table 1. systematic uncertainty was estimated at 150 K due to intrinsic scatter in the temperature correction from Frebel et al. (2013a). The corresponding systematic uncertainties for the remaining stellar parameters, , , and , were set to 0.3 dex, 0.2 km s-1, and 0.2 dex, respectively. These are added in quadrature to the statistical uncertainties in the slopes and standard errors of the spectroscopic balance to determine the total stellar parameter uncertainties (see Ji et al., 2020a, for details).
Name | [Fe/H] | |||||||
Leiptr-186 | 4640 | 152 | 1.00 | 0.31 | 2.51 | 0.24 | 0.21 | |
Leiptr-208 | 5020 | 157 | 1.90 | 0.31 | 1.85 | 0.24 | 0.21 | |
Leiptr-252 | 5360 | 163 | 3.60 | 0.31 | 1.31 | 0.25 | 0.33 | |
Leiptr-321 | 5095 | 160 | 2.40 | 0.31 | 1.95 | 0.27 | 0.22 | |
Leiptr-342 | 4650 | 153 | 1.00 | 0.30 | 2.32 | 0.25 | 0.21 |
@@
Overall, we detected or put upper limits on 29 species of 24 elements, where CH, CN, Al I, Sc II, V I, V II, Mn I, Co I, Sr II, Y II, Ba II, La II, and Eu II were synthesized and O I, Na I, Mg I, K I, Ca I, Ti I, Ti II, Cr I, Cr II, Fe I, Fe II, Ni I, Zn I, Sr I, Zr II, and Dy II were determined with equivalent widths. Mean abundances for each species are computed with a weighted average of individual absorption lines. For an element X, the weighted abundance is:
where each line has an inferred abundance , and the inverse variance weight is
where , , , and are the per-line abundance uncertainties from propagating the stellar parameter uncertainties and is the abundance uncertainty due to the equivalent width uncertainty. We additionally adopt a minimum systematic uncertainty per line, = 0.1, which acts as an uncertainty floor to account for unknowns in aspects like the model atmosphere and atomic data. All values are given in Table 3. We adopt the Asplund et al. (2009) solar abundances.
Abundance uncertainties are determined following a similar method to the one used in Ji et al. (2020b). The statistical uncertainty, , is the combination of the weighted standard deviation and the total standard error:
@@
The propagated stellar parameter uncertainty, , is:
@@
Each of these terms is treated as the total uncertainty of the element X due to the corresponding stellar parameter, e.g.:
The total uncertainty for [X/H] is the quadrature sum:
@@
We use [Fe I/H] to calculate [X/Fe]. The uncertainty of [X/Fe] needs to account for statistical uncertainties for both X and Fe, in addition to correlated abundance uncertainties from stellar parameters:
@@
We present the start of a list of line measurements for the Leiptr candidate stars in Table 3 which provides each line’s wavelength, excitation potential, oscillator strength, equivalent width, and full width at half maximum. The full table is available in the arXiv source as a CSV file. The final abundances for our five stars are given in Tables 3-4.2. We record element, number of lines used, and log , as well as [X/H] and [X/Fe] and their respective uncertainties, and , for each star. gives the non-LTE (NLTE) abundances for sodium, aluminum, potassium, and manganese (see Sections 4.4 and 4.5).

### v2

@@ Abstract
Over the course of their lifetimes, stellar systems can be torn apart by the gravitational potential of massive galaxies. Stellar streams form during the intermediate stage of tidal disruption where they preserve their coherence in phase-space before mixing into a stellar halo (Helmi & White, 1999). Current models of galaxy formation predict that there should be a large population of globular cluster and dwarf galaxy stellar streams in the Milky Way’s stellar halo as a result of past accretion events (Helmi, 2020). These systems provide observational evidence to evaluate theoretical models of hierarchical structure formation (e.g., Johnston, 1998; Helmi & White, 1999; Freeman & Bland-Hawthorn, 2002; Bullock & Johnston, 2005) and quantify the amount of dark matter substructure (e.g., Ibata et al., 2002; Johnston et al., 2002; Carlberg, 2009; Varghese et al., 2011; Erkal et al., 2016b). Large photometric surveys such as the Sloan Digital Sky Survey (York et al. 2000) and the Dark Energy Survey (Dark Energy Survey Collaboration et al. 2016) in combination with proper motions from Gaia data releases (Gaia Collaboration et al., 2016; Lindegren et al., 2021; Gaia Collaboration et al., 2023) have revealed over 140 streams (Grillmair & Carlin, 2016; Shipp et al., 2018; Mateu et al., 2018; Mateu, 2023; Ibata et al., 2024; Bonaca & Price-Whelan, 2025).
@@
It is often unclear whether a given stream’s progenitor is a dwarf galaxy, a globular cluster, or possibly a superposition of both (e.g., Carlberg, 2018; Malhan et al., 2021). In some cases, streams are clearly identified with intact progenitors, such as the Sagittarius dwarf galaxy (Ibata et al., 2001; Majewski et al., 2003), the Antlia II and Crater II dwarf spheroidals (Ji et al., 2021; Coppi et al., 2024), and several globular clusters (e.g., Simpson et al., 2020; Hansen et al., 2020b; Gull et al., 2021). The majority of streams do not have associated intact progenitors. Instead, a stream’s morphology and kinematics can provide a strong prior on the progenitor (e.g., Erkal et al., 2016a; Shipp et al., 2018; Ibata et al., 2024). However, a definitive classification requires studying a stream’s chemical abundances (e.g., Li et al., 2019, 2022; Ji et al., 2020a; Martin et al., 2022).
@@
The majority of known streams are thought to have globular cluster progenitors, due to their thin widths and in many cases small or unresolved metallicity dispersions dex (Li et al., 2022; Ibata et al., 2024). These include the Phoenix (Wan et al., 2020; Casey et al., 2021) and C-19 streams (Martin et al., 2022; Yuan et al., 2022a) which are below the globular cluster metallicity “floor”; the dynamically perturbed GD-1 (Grillmair & Dionatos, 2006; Price-Whelan & Bonaca, 2018) and ATLAS-Aliqa Uma streams (Shipp et al., 2018; Li et al., 2021); disrupted globular clusters embedded within the Indus, Jhelum, and Wukong/LMS-1 dwarf galaxy streams (Hansen et al., 2021; Awad et al., 2024; Limberg et al., 2024); and the 300S stream, which was originally thought to be a dwarf galaxy stream (Simon et al., 2011; Frebel et al., 2013b; Fu et al., 2018) but has since been chemically confirmed to be an accreted globular cluster stream (Li et al., 2022; Usman et al., 2024).
@@
Less common by number are the dwarf galaxy streams. Many of the first discovered streams, such as the Helmi stream (Helmi et al., 1999; Roederer et al., 2010; Limberg et al., 2021b; Matsuno et al., 2022b) and Sagittarius stream (Ibata et al., 2001; Majewski et al., 2003), were easily identified as dwarf galaxies as these are well-populated streams from relatively massive () dwarf galaxy progenitors. More recently, several lower-mass dwarf galaxy streams have also been identified. Based on their chemical patterns, the Indus and Jhelum streams have progenitor mass (Hansen et al., 2021); the Orphan-Chenab stream (Belokurov et al., 2007; Casey et al., 2013; Shipp et al., 2018; Koposov et al., 2019; Erkal et al., 2019) has progenitor mass (Koposov et al., 2019; Hawkins et al., 2023); the Wukong/LMS-1 stream has progenitor mass (Naidu et al., 2020; Yuan et al., 2020; Limberg et al., 2024); the Cetus/Palca stream has progenitor mass (Newberg et al., 2009; Yuan et al., 2022b; Sitnova et al., 2024); and the Elqui stream has progenitor mass (Shipp et al., 2018; Ji et al., 2020a; Li et al., 2022). This does not include more disrupted and phase-mixed dwarf galaxies like Gaia-Sausage/Enceladus (GSE, Belokurov et al., 2018; Haywood et al., 2018; Helmi et al., 2018; Carrillo et al., 2022), Sequoia (Myeong et al., 2019; Matsuno et al., 2019), Thamnos (Koppelman et al., 2019), and others (Naidu et al., 2020; Helmi, 2020; Deason & Belokurov, 2024).
@@
Current predictions for globular cluster streams are that the Milky Way has globular cluster stellar streams that have yet to be fully phase-mixed. These are predicted by mock stellar streams from the orbital histories of progenitors generated by hierarchical models of globular cluster formation (Pearson et al., 2024). There are also expected to be many hundreds of dwarf galaxies that have been tidally disrupted around the Milky Way. A number of strategies have been developed for this over time, such as semi-analytic methods for determining phase-space populations of tidal debris, simulations of satellite galaxy disruption, and clustering stellar dynamics in 4D phase space (e.g., Johnston, 1998; Helmi & White, 1999; Santistevan et al., 2020; Brauer et al., 2022; Gandhi et al., 2024). However, it is harder to identify the dwarf galaxy streams. Current stream searches, especially those relying on Gaia parallaxes (e.g., Ibata et al., 2019, 2024; Dodd et al., 2023), mostly search at nearby galactocentric radii, where globular cluster streams are likely more numerous than dwarf galaxy streams (Johnston et al., 2008; Gómez et al., 2013; Pearson et al., 2024; Bonaca & Price-Whelan, 2025). Dwarf galaxy stars have larger velocity dispersions and metallicity spreads, so at fixed stellar mass they are harder to find compared to globular clusters that are more spatially compact and at single metallicities. Finally, the vast majority of the tidally disrupted dwarf galaxies are very low-mass, with , i.e. ultra-faint dwarf galaxies (e.g., Weisz & Boylan-Kolchin, 2017; Simon, 2019; Brauer et al., 2019; Santistevan et al., 2020; Gandhi et al., 2024), so they should be even harder to identify. Obtaining a full census of the Milky Way’s accretion history thus requires extending stream discoveries down to this regime of ultra-faint dwarf galaxies. Since these ultra-faint dwarf galaxies are also the most dark matter-dominated galaxies that live in the lowest-mass star-forming dark matter halos, understanding the number of these streams in the Milky Way could contribute to understanding reionization and the small scale structure of dark matter (e.g., Bullock & Boylan-Kolchin, 2017; Jethwa et al., 2018; Simon, 2019; Nadler et al., 2021).
@@
In recent years, much progress has been made in finding streams from the lowest-mass dwarf galaxies. The first clear stream around an intact ultra-faint dwarf galaxy was found around the Tucana III system (Drlica-Wagner et al., 2015; Shipp et al., 2018), which has been spectroscopically confirmed (Hansen et al., 2017; Li et al., 2018). More recently, a tidal stream likely emanating from the Hydrus I dwarf galaxy was found (Ibata et al., 2024). Several individual stars with chemistry similar to ultra-faint dwarf galaxy stars have also been found throughout the Milky Way halo (e.g., Casey & Schlaufman, 2017; Roederer, 2017). Motivated by the discovery of the highly -process (rapid neutron-capture process) enhanced ultra-faint dwarf Reticulum II (Ji et al., 2016a; Roederer et al., 2016), several stellar dynamical groups within the Milky Way that are highly enhanced in -process elements have also been suggested to originate from ultra-faint dwarf galaxies (e.g., Roederer et al., 2018; Lee et al., 2019; Limberg et al., 2021a; Gudin et al., 2021), but the probability that the stars in these groups all come from individual ultra-faint dwarf galaxies is statistically unlikely (Brauer et al., 2022). Thus far, there is no progenitor-free bona-fide stellar stream that can be cleanly associated with an ultra-faint dwarf galaxy. The only claimed candidate, Specter (Chandra et al., 2022), has a high mean metallicity that suggests we are only seeing part of the stream from a much higher-mass progenitor dwarf galaxy.
@@
Here we investigate the progenitor of the Leiptr stellar stream, which was previously thought to be a tidally disrupted globular cluster. Leiptr was discovered among a group of inner galaxy streams in the Gaia DR2 catalog by Ibata et al. (2019) using the STREAMFINDER algorithm. They identified candidates from streams with distinct and coherent sky positions, proper motions, and parallaxes that were clearly distinguishable from the Galactic disk or bulge. They determined Leiptr was 48∘ long and located toward the Galactic anticenter with a heliocentric distance of kpc, pericenter distance of kpc, apocenter distance of kpc, and strong retrograde motion of = km s-1 kpc where is the vertical component of angular momentum in a Galactocentric Cartesian frame. Malhan et al. (2022) updated these orbital parameters, estimating the pericenter to be 12.3 0.1 kpc and the apocenter to be 45.1 0.2 kpc, with a resulting eccentricity of 0.57. Leiptr has the highest of stellar streams in their sample, next to Gaia-12. Bonaca et al. (2021) identified Leiptr, Gjöll, GD-1, Phlegethon, Ylgr, and Wambelong as potential members of a highly aligned “plane of streams” that may share a common origin. They suggested the retrograde streams are most likely globular clusters from the Sequoia/Arjuna dwarf galaxy (Myeong et al., 2019; Naidu et al., 2020) based on photometry and clustering in orbital poles; Gjöll in particular has been shown to be connected to the NGC 3201 globular cluster (Hansen et al., 2020b; Riley & Strigari, 2020). Bonaca & Price-Whelan (2025) independently determined that Leiptr has a mean heliocentric distance of 7.4 kpc, mean Galactic radius of 13.8 kpc, angular length of 73∘, angular width of 0.5∘ (74 pc), and a stellar mass of . However, there have been no previous chemical studies of Leiptr.
We present the chemical abundances of four member stars in the Leiptr stream and one non-member using high-resolution spectroscopy. We find that Leiptr is likely the first spectroscopically confirmed ultra-faint dwarf galaxy stream without an intact progenitor. Section 2 explains the observations. Section 3 describes the abundance analysis. Section 4 compares the chemical abundances of Leiptr to those of halo, dwarf galaxy, and globular cluster stars. We discuss what these results suggest about Leiptr’s chemical signature and progenitor in Section 5 and conclude in Section 6.
@@ 2 Observations
Of 67 original Leiptr candidate stars from the Gaia DR2 STREAMFINDER catalog (Ibata et al., 2019), we selected five of the brightest candidate stars to observe, giving them arbitrary ID numbers. We obtained spectra of these stars on November 2, 2022 using the 6.5-meter Magellan Clay Telescope’s Magellan Inamori Kyocera Echelle (MIKE; Bernstein et al. 2003) spectrograph at Las Campanas Observatory in Chile. Figure 1 shows the declinations and radial velocities against right ascensions of these stars compared to the other Leiptr candidates from the Gaia DR3 STREAMFINDER catalog in Ibata et al. (2024). The arrows represent reflex-corrected proper motions. One of these five stars, Leiptr-252, is no longer a kinematic Leiptr member star in the Gaia DR3 STREAMFINDER catalog (Ibata et al., 2024), which will also be apparent from its chemistry. This non-member star is shown as a red X in Figure 1 and all subsequent figures.
@@ 3 Abundance Analysis
The chemical abundance analysis was performed using a custom analysis code LESSPayne (Ji et al., 2025), generally following Ji et al. (2020a). As described in Limberg et al. (2024), LESSPayne is a combination of the PayneEchelle (Ting et al., 2019) and SMHR (Casey, 2014) programs. First, we ran a full spectrum fit between Å using a neural network emulator of high-resolution spectra. We did not use the stellar parameters from this fit, but instead used the best-fit spectrum to locate and mask regions where absorption lines were expected to occur and initialize an SMHR analysis. We manually inspected and corrected the normalization for every spectral order before stitching them together. We then fit Gaussian profiles to determine equivalent widths using the line list from Ji et al. (2020a), which is based on atomic data from linemake (Placco et al., 2021), mostly keeping to lines with 4000 Å 7000 Å. Some stronger lines required Voigt profile fits. We also fit some lines using spectral synthesis with the same line lists as Ji et al. (2020a).
@@
We determined chemical abundances using the 1D local thermodynamic equilibrium (LTE) radiative transfer code MOOG (Sneden, 1973) and the ATLAS model atmospheres (Castelli & Kurucz, 2003). Stellar parameters were determined spectroscopically following Frebel et al. (2013a), due to uncertain distances to the Leiptr stars. Effective temperature was derived by balancing Fe I abundance against excitation potential, surface gravity by balancing Fe I and II abundances, and microturbulence by balancing Fe II abundance against reduced equivalent width. Next, we applied the correction to a photometric temperature scale from Frebel et al. (2013a) and redetermined the other stellar parameters. We estimate the microturbulence using Fe II lines because of the photometric correction: using Fe I lines computed in LTE with photometric temperatures results in artificially high microturbulence (Ji et al., 2020a). This is possible as we used Fe II lines per star. The corrected effective temperature (), surface gravity (log g), microturbulence (), metallicity ([Fe/H]), and their associated uncertainties (, , , and , respectively) are given in Table 3. systematic uncertainty was estimated at 150 K due to intrinsic scatter in the temperature correction from Frebel et al. (2013a). The corresponding systematic uncertainties for the remaining stellar parameters, , , and , were set to 0.3 dex, 0.2 km s-1, and 0.2 dex, respectively. These are added in quadrature to the statistical uncertainties in the slopes and standard errors of the spectroscopic balance to determine the total stellar parameter uncertainties (see Ji et al., 2020a, for details).
cccccccccc
[…]
Leiptr-208 5020 157 5126 1.90 0.31 1.85 0.24 0.21
Leiptr-252 5360 163 5329 3.60 0.31 1.31 0.25 0.33
Leiptr-321 5095 160 5059 2.40 0.31 1.95 0.27 0.22
Leiptr-342 4650 153 4513 1.00 0.30 2.32 0.25 0.21
[…]
is spectroscopic effective temperature with the correction from Frebel et al. (2013a), log g is surface gravity, and is microturbulence.
We also include the photometric temperature , as calculated using Mucciarelli et al. (2017) color-temperature relations for Gaia BP-RP colors.
The values represent the stellar parameters’ respective uncertainties, which are a quadrature sum of the statistical uncertainty calculated from the iron lines and adopted systematic uncertainties for Teff, , , and of 150 K, 0.3 dex, 0.2 km s-1, and 0.2 dex, respectively.
@@
Overall, we detected or put upper limits on 29 species of 24 elements, where CH, CN, Al I, Sc II, V I, V II, Mn I, Co I, Sr II, Y II, Ba II, La II, and Eu II were synthesized and O I, Na I, Mg I, K I, Ca I, Ti I, Ti II, Cr I, Cr II, Fe I, Fe II, Ni I, Zn I, Sr I, Zr II, and Dy II were determined with equivalent widths. Mean abundances for each species are computed with a weighted average of individual absorption lines and compared to their respective solar abundances from Asplund et al. (2009).
For a given absorption line for element , we measure an abundance . The measured abundance is uncertain, due to uncertainties in the stellar parameters and the measurement itself. We therefore measure abundance uncertainties for each absorption line by independently propagating the stellar parameter uncertainties, denoted , , , and . Additionally, uncertainty in the measured equivalent width propagates into an abundance uncertainty denoted .
To measure the overall abundance for the star, we use a weighted average of measured abundances from all absorption lines. Weighting is introduced in order to de-prioritize measurements with high errors to most accurately represent the measured abundance of the star. The weighted abundance is calculated as:
The weight for a measured line reflects the total uncertainty of a measurement by combining all of the measured uncertainties described above. The weight is calculated as the inverse of the quadrature sum of each contributing source of uncertainty:
We also calculate the total standard error, , which represents the contributed systematic and statistical uncertainties from each individual line measurement to the overall error of the element abundance, and is calculated as:
We set the per-line systematic error floor to account for unforeseen systematics in our stellar models and inferred abundance measurements.
We calculate the variance of our measurements, and continue to weight the contribution of each measurement, to appropriately estimate the systematic uncertainty in line-to-line scatter. The square of the weighted standard deviation is calculated as:
@@
We also calculate the contribution of the uncertainty relative to each stellar parameter. The propagated stellar parameter uncertainty, , is:
@@
Each of these terms is treated as the total uncertainty of the element due to the corresponding stellar parameter, e.g.:
Here, we neglect correlations between stellar parameters. Ji et al. (2020a) gave a formalism for how to propagate correlated stellar parameter uncertainties to abundances, but it required the assumption that abundances depend linearly on stellar parameters. We have unfortunately found that this assumption often breaks down, which can result in systematic biases in the resulting abundances. By ignoring the correlations, we find that we slightly overestimate the stellar parameter uncertainties, which we find preferable to obtaining biased abundances.
The total uncertainty for the abundance of element in the star, [X/H], combines all previous uncertainties. is the quadrature sum of the total standard error, the stellar parameter uncertainties, and the weighted standard deviation:
We include a minimum floor of on the weighted variance of measurements to account for elements with too few lines to measure a standard deviation. In our tables and figures, we also include the abundance uncertainty excluding this systematic uncertainty buffer and weighted variance:
We present both sets of uncertainties, one that gives our best estimate of a total uncertainty, and the other to estimate abundance precision.
@@
We use [Fe I/H] to calculate [X/Fe]. The uncertainty of [X/Fe] needs to account for statistical uncertainties for both X and Fe, in addition to correlated abundance uncertainties from stellar parameters and the weighted variance with corresponding uncertainty floors:
Again, we exclude the uncertainty floor to give an approximate error without a conservative error buffer:
@@
We present the start of a list of line measurements for the Leiptr candidate stars in Table 3 which provides each line’s wavelength, excitation potential, oscillator strength, equivalent width, and full width at half maximum. The full table is available in the arXiv source as a CSV file. The SMHR analysis files that include the spectra and line fits are available upon request to the first author. The final abundances for our five stars are given in Tables 1-4.2. We record element, number of lines used, and log , as well as [X/H] and [X/Fe] and their respective uncertainties, and , for each star. gives the non-LTE (NLTE) abundances for sodium, aluminum, potassium, and manganese (see Sections 4.4 and 4.5). We comment on the NLTE effects for all significantly impacted elements, ignoring smaller contributions that are not informative to this analysis.